# Human-In-the-Loop Software Development Agents

### v1

@@ V-C Stage 3: Practitioners’ Survey
To gather deeper insights on HULA quality and feedback, we conducted an online survey with practitioners who have used HULA to generate code (i.e., reach to the refine code step) at least once. The survey is design to capture the performance of HULA in the internal deployment and to gather the user perceptions (e.g., benefits, challenges, and areas of improvement). We conduct the pilot survey with five software engineers to validate and refine the questions prior to the full launch. In the full launch, we sent a survey invitation directly to the practitioners via Slack communication.
@@ VI Results
Table III shows that HULA can correctly identify all the files (i.e., a Perfect F1 score) for 84% of the issues in SWE-bench. The code generated by HULA achieved a similarity score of 45% compared to human-written code and passed the unit tests for 31% of the issues. This performance is comparable to SWE-agent Claude, which ranks 6th on the SWE-bench leaderboard. These results indicate that HULA performs reasonably well on the SWE-bench dataset.
Table III shows that HULA achieved a perfect F1 of 15% and a similarity score of 30% compared to human-written code in our internal dataset. The lower performance could be due to various factors. One possible factor is the diversity of the data. The SWE-bench dataset is limited to the Python programming language, whereas our internal dataset includes more than 10 programming languages and over 20 file types. Moreover, the SWE-bench dataset was collected from 12 open-source software repositories, while our internal dataset was sourced from 94 software repositories. This increased diversity in programming languages and repositories may challenge the LLM agent’s generalizability, leading to suboptimal performance.
@@
Figure 4 shows that out of 433 issues that had the coding plans approved, HULA successfully generated code changes for 376 issues, achieving a code generation rate of 87%. Of these 376 issues, the HULA-generated code for 95 issues was then raised as PRs for review by other practitioners, achieving a raised PR rate of 25%. Eventually, 56 PRs out of the 95 PRs that contained the HULA-generated code were merged, acheiving a merged PR rate of 59%.
@@
Finding 2. We find that during the planning stage, practitioners approved the generated coding plans for 433 out of 527 plan generated issues, achieving a plan approval rate of 82%. During the coding stage, pull requests were created for 95 out of 376 issues by HULA, achieving a raised PR rate of 25%. Of these, 56 pull requests were successfully merged, leading to a merged PR rate of 59%.
@@
Figure 6 presents the survey responses to our Likert-scale questions regarding the performance of HULA during the planning and coding stages. Based on 109 survey responses, we generally observe low user disagreements in the planning stage. Particularly, a total of 82% and 76% of participants either agreed or responded neutrally that the identified files were relevant to the issue description (Q1) and aligned with how they would have approached the task (Q2), respectively. Considering the generated plans, 69% and 64% of the participants either agreed or responded neutrally that the change plans aligned with the issue description (Q3) and aligned with how they would have approached the task (Q4). These results provide further insight into HULA’s strong capability to identify relevant files and plan to mimic user behavior.
Figure 6 shows that 83% of the participants either agreed or responded neutrally that the generated code could be easily understood and modified to solve the issue (Q5). However, an adjustment is likely needed after the first generation. Without human involvement, 50% of participants either agreed or responded neutrally that the generated code solved their task (Q6). Yet, 54% and 67% of them disagreed that the code contains no defects and fulfilled non-functional requirements (Q7) and completely solved the task without human involvement (Q8), respectively. These results suggest that the participants were satisfied with the generated code while they could adjust the code to match their needs.
@@
Figure 9 presents the analysis of the open-ended question regarding the areas of improvement for HULA from participants’ perspective (Q11). Regarding the improvement in the UX/UI workflow (17 responses), participants highlighted that they prefer the flexibility to move between planning and coding stages to provide follow-up feedback (i.e., the revisiting workflow), particularly when handling partially incomplete changes. Additionally, participants suggested to automatically enhance the input context with documentations (e.g., type definitions), best practices (e.g., design patterns), and other historical data (e.g., previous solutions) (14 responses). Other suggestions include providing input guidance for users, enhancing code validation with selectable options, and extending capabilities to generate test cases for corresponding changes.
Finding 3. Most participants agreed that the coding plan was accurate and the generated code was easy to read and modify, which helped reduce their initial development time and effort. However, human involvement is still needed to ensure the code completely addresses the task. In addition, participants acknowledged that HULA promotes good documentation practice, but this requires more effort to provide detailed issue descriptions.
@@ VII Lessons Learned and Future Directions
Lesson Learned 1: The performance of LLM-based software development agent heavily relies on a detailed input description, but what key information is needed? In RQ1, we found that HULA achieves a lower accuracy on the internal dataset when compared to the SWE-bench dataset. One of the possible reasons is related to the nature of how a task is written. For the SWE-bench dataset, issues typically have a detailed description along with key information e.g., module names or code snippets. However, for Agile-driven organizations like Atlassian, practitioners tend to collaborate closely through informal knowledge transfer (e.g., meetings, discussion threads, or chats). Therefore, Agile-driven JIRA issues may contain brief information like a user story or a one-liner description instead of detailed documentation (as observed in our internal dataset).
@@
On the other hand, RQ2 indicates that practitioners agreement with the generation is high when they can engage in the process by reviewing and enriching the issue descriptions with additional information. Although in RQ3, a few participants in the survey acknowledged that HULA’s workflow could promote good documentation practice, it requires more effort to provide detailed issue descriptions. To address this challenge, participants requested guidance on the needed types of input information. Yet, little is known what key information is needed. Future research directions should investigate: What types of information must be presented for LLM-based software development agents?, How to automatically augment such information to LLM-based software development agents?, What is the best augmentation approach for LLM-based software development agents?
@@
Lesson Learned 2: Evaluating functional correctness should go beyond passing unit test cases. Functional correctness is the ultimate goal of any LLM-based software development agents (i.e., to generate code that is functionally correct). During our offline evaluation, we followed the state-of-the-art approaches [31, 32, 33] to evaluate the code functionality based on the percentage of the passing unit test cases (same as SWE-bench [24]). However, our RQ3 shows that HULA may not always generate functionally correct code despite incorporating human and tools feedback.
@@
Passing unit test cases should not be the sole goal of evaluating code functionality due to the following reasons. First, passing unit test cases requires practitioners to prepare both test cases and the test environment (e.g., build and configuration files) to execute test cases against the code changes for a specific programming language for every pull request. However, this approach is highly cost-intensive and intricate, particularly in our enterprise context, due to the complexity of the task and software dependencies. Secondly, the binary nature of the test results (i.e., pass or fail) lacks the granularity required for the team to assess how closely the generated code aligns with the ground truth. Third, test cases only show the presence of the software defects, but not their absence. Therefore, the generated code that passes unit test cases does not guarantee that the code is functionally correct. To address this challenge, future research directions should explore: How to generate explanations or descriptions from test execution that are more beneficial to practitioners and LLM-based agents when test cases are available?, What is an alternative approach to evaluate code functionality when test cases are not available?, How to go beyond the traditional evaluation of functionality correctness (e.g., readability, technical debt, maintainability, etc)?
@@ VIII Threat to validity
Threats to construct validity. Measuring code similarity is complex and nuanced. Code similarity can be measured based on various dimensions, e.g., semantic similarity, syntactic similarity, functional similarity. To ensure that LLM-as-a-Judge for code similarity is highly similar to human judgement, we developed prompts that capture various concepts of code similarity. To validate if the code similarity score is aligned with human judgement, we randomly selected 203 Jira issues. We asked software engineers to provide a similarity score, ranging from 1 (no similarity), 2 (weak similarity), 3 (high similarity), to 4 (highly similarity). Then, we computed the correlation efficient between the human score and LLM-as-a-Judge score. We observed a high correlation coefficient of 0.7, demonstrating the high degree of alignment between human judgement and LLM-as-a-judge.
Threats to external validity. Our HULA framework is generally applicable to other coding platforms and other LLMs. However, for this paper, the findings are limited to the context of Atlassian Jira that are internally used by Atlassian software engineers, which may not be generalized to other contexts. Therefore, the application of this framework in other coding platforms will prove fruitful.
Threats to internal validity. We designed our HULA framework to be architecture-agnostic, meaning that other LLMs are applicable to this framework. However, the experiment and findings of this paper rely on the use of GPT-4 as a backbone LLM for our HULA framework. Based on the SWE-Bench Leaderboard, we find that our performance is comparable to the 6th-ranked agent, highlighting the competitive performance of our backbone LLM. It is possible that other LLMs that are larger with higher number of parameters may achieve higher performance. Nevertheless, the key goal of this paper is not to investigate which LLMs perform best, but instead to investigate the effectiveness of HULA framework in the real-world deployment. Therefore, the investigation of the best-performing LLMs is left for future work.
Lastly, although we may not be able to disclose all aspects of the implementation and experiment design due to the industrial confidentiality reasons, we are confident that the successful development and integration of HULA framework into Atlassian Jira is a great achievement, paving the way forward the future of AI-powered software development tools. In addition, our experience, findings, lessons learned, and future research directions will be beneficial for other practitioners and software engineering researchers.
@@ IX Conclusion
In this paper, we introduce a Human-in-the-Loop LLM-based Software Development Agents Framework (HULA). We design the user interface, implement, and deploy of LLM-based software development agents into Atlassian JIRA for internal uses. Through a multi-stage evaluation of HULA involving real-world deployment and user surveys with Atlassian practitioners, we conclude that (1) the detail of input can highly effect the performance of HULA; (2) incorporating human feedback into HULA can enhance the input context and be beneficial in practice; (3) practitioners perceive that HULA can help minimize the overall development time and effort but code quality concerns are raised in some cases.
X Disclaimer

### v2

@@ V-C Stage 3: Practitioners’ Survey
To gather deeper insights on HULA quality and feedback, we conducted an online survey with practitioners who have used HULA to generate code (i.e., reach the refine code step) at least once. The survey is designed to capture the performance of HULA in the internal deployment and to gather the user perceptions (e.g., benefits, challenges, and areas of improvement). We conduct the pilot survey with five software engineers to validate and refine the questions prior to the full launch. In the full launch, we sent a survey invitation directly to the practitioners via Slack communication.
@@ VI Results
Table III shows that HULA can correctly identify all the files (i.e., Perfect File Localization) for 84% of the issues in SWE-bench. The code generated by HULA achieved a similarity score of 45% compared to human-written code and passed the unit tests for 31% of the issues. This performance is comparable to SWE-agent Claude, which ranks 6th on the leaderboard as of 26 September 2024. These results indicate that HULA performs reasonably well on SWE-bench.
Table III shows that HULA achieved a perfect file localization on 15% of the issues and a high similarity score compared to human-written code on 30% of the issues in our internal dataset. The lower performance could be due to various factors. One possible factor is the diversity of the data. The SWE-bench dataset is limited to the Python programming language, whereas our internal dataset includes more than 10 programming languages and over 20 file types. Moreover, the SWE-bench dataset was collected from 12 open-source software repositories, while our internal dataset was sourced from 94 software repositories. This increased diversity in programming languages and repositories may challenge the LLM agent’s generalizability, leading to suboptimal performance.
@@
Figure 4 shows that out of 433 issues that had the coding plans approved, HULA successfully generated code changes for 376 issues, achieving a code generation rate of 87%. Of these 376 issues, the HULA-generated code for 95 issues was then raised as PRs for review by other practitioners, achieving a raised PR rate of 25%. Eventually, 56 PRs out of the 95 PRs that contained the HULA-generated code were merged, achieving a merged PR rate of 59%.
@@
Finding 2. We find that during the planning stage, practitioners approved the generated coding plans for 433 out of 527 plan-generated issues, achieving a plan approval rate of 82%. During the coding stage, pull requests were created for 95 out of 376 code-generated issues, achieving a raised PR rate of 25%. Of these, 56 pull requests were successfully merged, leading to a merged PR rate of 59%.
@@
Figure 6 presents the survey responses to our Likert-scale questions regarding the performance of HULA during the planning and coding stages. Based on 109 survey responses, we generally observe low user disagreements in the planning stage. Particularly, a total of 71% and 62% of participants agreed that the identified files were relevant to the issue description (Q1) and aligned with how they would have approached the task (Q2). Considering the generated plans, 41% and 36% of the participants agreed that the change plans aligned with the issue description (Q3) and aligned with how they would have approached the task (Q4). These results provide further insight into HULA’s strong capability to identify relevant files and plan to mimic user behaviour.
Figure 6 shows that 61% of the participants agreed that the generated code could be easily understood and modified to solve the issue (Q5). However, an adjustment is likely needed after the first generation. Without human involvement, 33% of the participants agreed that the generated code solved their task (Q6). Yet, 54% and 67% of them disagreed that the code contains no defects and fulfilled non-functional requirements (Q7) and completely solved the task without human involvement (Q8), respectively. These results suggest that HULA still faces challenges in generating the perfect code, however, the generated code is easy to modify to match participants’ needs.
@@
Figure 9 presents the analysis of the open-ended question regarding the areas of improvement for HULA from the participants’ perspective (Q11). Regarding the improvement in the UX/UI workflow (17 responses), participants highlighted that they prefer the flexibility to move between planning and coding stages to provide follow-up feedback (i.e., the revisiting workflow), particularly when handling partially incomplete changes. Additionally, participants suggested automatically enhancing the input context with documentation (e.g., type definitions), best practices (e.g., design patterns), and other historical data (e.g., previous solutions) (14 responses). Other suggestions include providing input guidance for users, enhancing code validation with selectable options, and extending capabilities to generate test cases for corresponding changes.
Finding 3. 41% of participants agreed that the generated plan by HULA aligned accurately with the issue description. While HULA still requires human involvement to ensure the generated code completely addresses the task, 61% of the participants agreed that the generated code was easy to read and modify, which helped reduce their initial development time and effort. Additionally, participants acknowledged that HULA promotes good documentation practice, though this requires more effort to provide detailed issue descriptions.
@@ VII Lessons Learned and Future Directions
Lesson Learned 1: The performance of an LLM-based software development agent heavily relies on a detailed input description, but what key information is needed? In RQ1, we found that HULA achieves a lower accuracy on the internal dataset when compared to the SWE-bench dataset. One of the possible reasons is related to the nature of how a task is written. For the SWE-bench dataset, issues typically have a detailed description along with key information e.g., module names or code snippets. However, for Agile-driven organizations like Atlassian, practitioners tend to collaborate closely through informal knowledge transfer (e.g., meetings, discussion threads, or chats). Therefore, Agile-driven JIRA issues may contain brief information like a user story or a one-liner description instead of detailed documentation (as observed in our internal dataset).
@@
On the other hand, RQ2 indicates that practitioners’ agreement with the generation is high when they can engage in the process by reviewing and enriching the issue descriptions with additional information. Although in RQ3, a few participants in the survey acknowledged that HULA’s workflow could promote good documentation practice, it requires more effort to provide detailed issue descriptions. To address this challenge, participants requested guidance on the needed types of input information. Yet, little is known what key information is needed. Future research directions should investigate: What types of information must be presented for LLM-based software development agents? How to automatically augment such information to LLM-based software development agents? What is the best augmentation approach for LLM-based software development agents?
@@
Lesson Learned 2: Evaluating functional correctness should go beyond passing unit test cases. Functional correctness is the ultimate goal of any LLM-based software development agent (i.e., to generate code that is functionally correct). During our offline evaluation, we followed the state-of-the-art approaches [39, 40, 41] to evaluate the code functionality based on the percentage of the passing unit test cases (same as SWE-bench [32]). However, our RQ3 shows that HULA may not always generate functionally correct code despite incorporating human and tool feedback.
@@
Passing unit test cases should not be the sole goal of evaluating code functionality due to the following reasons. First, passing unit test cases requires practitioners to prepare both test cases and the test environment (e.g., build and configuration files) to execute test cases against the code changes for a specific programming language for every pull request. However, this approach is highly cost-intensive and intricate, particularly in our enterprise context, due to the complexity of the task and software dependencies. Secondly, the binary nature of the test results (i.e., pass or fail) lacks the granularity required for the team to assess how closely the generated code aligns with the ground truth. Third, test cases only show the presence of the software defects, but not their absence. Therefore, the generated code that passes unit test cases does not guarantee that the code is functionally correct. To address this challenge, future research directions should explore: How to generate explanations or descriptions from test execution that are more beneficial to practitioners and LLM-based agents when test cases are available? What is an alternative approach to evaluating code functionality when test cases are not available? How to go beyond the traditional evaluation of functionality correctness (e.g., readability, technical debt, maintainability, etc)?
@@ VIII Threat to validity
Threats to construct validity. Measuring code similarity is complex and nuanced. Code similarity can be measured based on various dimensions, e.g., semantic similarity, syntactic similarity, and functional similarity. To ensure that LLM-as-a-Judge for code similarity is highly similar to human judgement, we developed prompts that capture various concepts of code similarity. To validate if the code similarity score is aligned with human judgement, we randomly selected 203 Jira issues. We asked software engineers to provide a similarity score, ranging from 1 (no similarity), 2 (weak similarity), 3 (high similarity), to 4 (highly similarity). Then, we computed the correlation efficient between the human score and LLM-as-a-Judge score. We observed a high correlation coefficient of 0.7, demonstrating the high degree of alignment between human judgement and LLM-as-a-judge.
Threats to external validity. Our HULA framework is generally applicable to other coding platforms and other LLMs. However, for this paper, the findings are limited to the context of Atlassian Jira which is internally used by Atlassian software engineers and may not be generalized to other contexts. Therefore, the application of this framework in other coding platforms will prove fruitful.
Threats to internal validity. We designed our HULA framework to be architecture-agnostic, meaning that other LLMs are applicable to this framework. However, the experiment and findings of this paper rely on the use of GPT-4 as a backbone LLM for our HULA framework. Based on the SWE-Bench Leaderboard, we find that our performance is comparable to the 6th-ranked agent as of 26 September 2024, highlighting the competitive performance of our backbone LLM. It is possible that other LLMs that are larger with a higher number of parameters may achieve higher performance. Nevertheless, the key goal of this paper is not to investigate which LLMs perform best, but instead to investigate the effectiveness of the HULA framework in real-world deployment. Therefore, the investigation of the best-performing LLMs is left for future work.
Lastly, although we may not be able to disclose all aspects of the implementation and experiment design due to industrial confidentiality reasons, we are confident that the successful development and integration of the HULA framework into Atlassian Jira is a great achievement, paving the way forward the future of AI-powered software development tools. In addition, our experience, findings, lessons learned, and future research directions will be beneficial for other practitioners and software engineering researchers.
@@ IX Conclusion
In this paper, we introduce a Human-in-the-Loop LLM-based Software Development Agents Framework (HULA). We design the user interface, implement the agents, and deploy the framework into Atlassian JIRA for internal uses. Through a multi-stage evaluation of HULA involving real-world deployment and user surveys with Atlassian practitioners, we conclude that (1) the detail of input can highly affect the performance of HULA; (2) incorporating human feedback into HULA can enhance the input context and be beneficial in practice; (3) practitioners perceive that HULA can help minimize the overall development time and effort but code quality remains a concern in some cases. While the generated plan is perceived as accurate, HULA still faces challenges in generating perfect code without human input. Hence, future research is encouraged to enhance the quality of LLM-generated code [42, 43].
X Acknowledgments
This work was carried out through a collaboration between Atlassian, Monash University, and the University of Melbourne. The authors sincerely thank the universities and Atlassian’s ADO AI team, with special appreciation to the engineering (Acrabat), data science, product, and design teams, for their invaluable support and contributions.
XI Disclaimer